# matrix,arrows,decorations.pathmorphing

### v2

@@ 3.4. The fixed point set of an antiholomorphic involution in the procongruence moduli stack
For $g(S)\leq 2$, the closed strata of the boundary of the fixed point locus $\overline{{\mathbb{M}}}(S)^{\iota}$ of codimension $k+1$ are parameterized by the fixed point set ${\check{C}}(S)_{k}^{\iota}$, for all $k\geq 0$. Moreover, ${\check{C}}(S)_{k}^{\iota}$ is the closure of the fixed point set $C(S)_{k}^{\iota}$ in the profinite set ${\check{C}}(S)_{k}$.
@@
From the proof of Proposition 3.5, it follows that the fixed point locus ${{\mathbb{M}}}(S)^{\iota}$ is the closure of the image of the fixed point locus ${\mathcal{T}}(S)^{\iota}$ in ${{\mathbb{M}}}(S)$ and that the fixed point locus $\overline{{\mathbb{M}}}(S)^{\iota}$ is the closure of the image of the fixed point locus $\overline{{\mathcal{T}}}(S)^{\iota}$ in $\overline{{\mathbb{M}}}(S)$. In particular, the boundary $\partial\overline{{\mathbb{M}}}(S)^{\iota}=\overline{{\mathbb{M}}}(S)^{\iota}\smallsetminus{{\mathbb{M}}}(S)^{\iota}$ is the closure of the boundary $\partial\overline{{\mathcal{T}}}(S)^{\iota}$ in the boundary $\partial\overline{{\mathbb{M}}}(S)=\overline{{\mathbb{M}}}(S)\smallsetminus{{\mathbb{M}}}(S)$. By Proposition 3.2, both claims of the proposition then follow. ∎
@@ 4.1. Preliminary lemmas
Before we proceed to the proof of Theorem 1.1, we need to prove first a series of lemmas. A subgroup of ${\check{\Gamma}}^{\pm}(S)$ is $\mathbb{I}$-characteristic if it is preserved by all elements of $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$.
@@
For all $g(S)\geq 0$, $\mathrm{P}{\check{\Gamma}}(S)$ is an $\mathbb{I}$-characteristic subgroup of $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$ and $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$ is an $\mathbb{I}$-characteristic subgroup of ${\check{\Gamma}}^{\pm}(S)$.
@@
The pure mapping class group $\mathrm{P}\Gamma(S)$ is the subgroup of the mapping class group ${\Gamma}(S)$ generated by Dehn twists. Hence, $\mathrm{P}{\check{\Gamma}}(S)$ is an $\mathbb{I}$-characteristic subgroup of both ${\check{\Gamma}}(S)$ and $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$. In order to prove that $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$ is an $\mathbb{I}$-characteristic subgroup of ${\check{\Gamma}}^{\pm}(S)$, it is enough to show that ${\check{\Gamma}}(S)$ is an $\mathbb{I}$-characteristic subgroup of ${\check{\Gamma}}^{\pm}(S)$.
It is then enough to show that $\operatorname{Inn}({\check{\Gamma}}(S))$ is a normal subgroup of $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$. This will follow if we show that, for any essential simple closed curve ${\gamma}$ on $S$, there is a character:
$\chi_{\gamma}\colon\thinspace\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\to{\widehat{\mathbb{Z}}}^{\ast},$
whose kernel contains $\operatorname{Inn}({\check{\Gamma}}(S))$ but not $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))$.
There is a natural representation $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}({\check{C}}(S))$, which, by Theorem 5.5 in [6], preserves topological types. Therefore, for any $f\in\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$, there is a $x\in{\check{\Gamma}}(S)$ such that $\operatorname{inn}x\circ f$ preserves the procyclic subgroup ${\hat{\mathrm{I}}}_{\gamma}$ generated by the Dehn twist $\tau_{\gamma}$. If $y\in{\check{\Gamma}}(S)$ is another such element, then $xy^{-1}$ fixes ${\gamma}$ and there holds $\operatorname{inn}(xy^{-1})(\tau_{\gamma})=\tau_{\gamma}$, so that $\operatorname{inn}(xy^{-1})$ acts trivially on ${\hat{\mathrm{I}}}_{\gamma}$.
Therefore, assigning to $f\in\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$ the automorphism induced by $\operatorname{inn}x\circ f$ on the subgroup ${\hat{\mathrm{I}}}_{\gamma}$, defines a representation $\chi_{\gamma}\colon\thinspace\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}({\hat{\mathrm{I}}}_{\gamma})$ which contains $\operatorname{Inn}({\check{\Gamma}}(S))$ in its kernel. On the other hand, for an element $x$ of ${\Gamma}^{\pm}(S)$ which is not orientation preserving, there holds $\chi_{\gamma}(\operatorname{inn}x)(\tau_{\gamma})=\tau_{\gamma}^{-1}$, which implies $\ker\chi_{\gamma}=\operatorname{Inn}({\check{\Gamma}}(S))$. ∎
From Lemma 4.1, it follows that the elements of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ are compatible with the augmentation map $\mathrm{P}{\check{\Gamma}}^{\pm}(S)\to\mathbb{Z}/2$ and, in particular, preserve the set of antiholomorphic involutions. Moreover, by Proposition 3.4, for $g(S)\leq 2$, the sets of conjugacy classes of antiholomorphic involutions in $\mathrm{P}\Gamma^{\pm}(S)$ and $\mathrm{P}{\widehat{\Gamma}}^{\pm}(S)$ can be identified. In particular, we have:
@@
For $g(S)=0$, there is only one ${\widehat{\Gamma}}(S)^{\pm}$-conjugacy class of antiholomorphic involutions in $\mathrm{P}{\widehat{\Gamma}}(S)^{\pm}$ on which $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\widehat{\Gamma}}(S)^{\pm})$ acts naturally.
@@
Antiholomorphic involutions in $\mathrm{P}\Gamma(S)^{\pm}$ do not swap the punctures of $S$. Therefore, all the punctures lie in $\operatorname{Fix}(\iota)$, which, in particular, is not empty. Since $g(S)=0$, $\operatorname{Fix}(\iota)$ is also connected and separating. Hence, an antiholomorphic involution in $\mathrm{P}\Gamma(S)^{\pm}$ is determined by the cyclic order of the punctures on $\operatorname{Fix}(\iota)$, so that two of them are conjugated by an element of ${\Gamma}(S)^{\pm}$. The conclusion then follows from Proposition 3.4. ∎
@@
For $g(S)=0$, the fixed point set $C(S)_{0}^{\iota}$ of an antiholomorphic involution $\iota\in\mathrm{P}\Gamma(S)^{\pm}$ is finite and consists of isotopy classes of simple closed curves on $S$ which have between them geometric intersection either 00 or $2$.
@@
The complement $S\smallsetminus\operatorname{Fix}(\iota)$ is the disjoint union of two unpunctured discs Let then $D^{\prime}$ and $D^{\prime\prime}$ be closed subdiscs of $S$ such that $D^{\prime}\cap D^{\prime\prime}=\operatorname{Fix}(\iota)$. For a representative $\alpha$ of $\{[\alpha]\}\in C(S)_{0}^{\iota}$ such that $\alpha=\iota(\alpha)$, the intersection $\alpha\cap D^{\prime}$ is a disjoint union of arcs with boundary in $\operatorname{Fix}(\iota)$. If $\alpha^{\prime}$ is one of such arcs, then $\alpha^{\prime}\cup\iota(\alpha^{\prime})\subset\alpha$ is a simple closed curve (so that $\alpha^{\prime}\cup\iota(\alpha^{\prime})=\alpha$), which crosses transversally $\operatorname{Fix}(\iota)$ in the two boundary points of $\alpha^{\prime}$. The isotopy class of $\alpha$ is then determined by the partition which $\alpha^{\prime}$ induces on the set of punctures of $S$ (which all lie on $\operatorname{Fix}(\iota)$). This implies the first statement of the proposition.
For $\{[\alpha]\}\neq\{[\beta]\}\in C(S)_{0}^{\iota}$ such that $\beta=\iota(\beta)$, the simple closed curve $\beta$ is the union of the two arcs $\beta^{\prime}=\beta\cap D^{\prime}$ and $\beta^{\prime\prime}=\beta\cap D^{\prime\prime}$ and it is clear that the arcs $\alpha^{\prime}$ and $\beta^{\prime}$ have geometric intersection either 00 or $1$. The second statement of the proposition then follows as well. ∎
By Lemma 4.1, there are natural homomorphisms $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ and $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$. We have:

### v3

@@ 3.5. The fixed-point set of an antiholomorphic involution in the procongruence moduli stack
For $g(S)\leq 2$, the fixed-point locus $\overline{{\mathbb{M}}}(S)^{\iota}$ of an antiholomorphic involution $\iota\in{\Gamma}^{\pm}(S)\subset{\widehat{\Gamma}}^{\pm}(S)$ is irreducible and contains ${{\mathbb{M}}}(S)^{\iota}$ as an open dense subspace.
@@
By [33, Theorem 3.6], there is a natural bijection between the set of connected components of the real locus of $({\mathcal{M}}(S)^{\lambda},\iota)$ and conjugacy classes of involutions in ${\Gamma}^{\lambda}\cdot\langle\iota\rangle$.
By (i) of Proposition 3.5 and Shapiro’s lemma, ${\Gamma}^{\lambda}\cdot\langle\iota\rangle$ is also good for $g(S)\leq 2$. By [10, Corollary B, (i)] and the above remarks, there is then a bijective correspondence between the set of connected components of the real locus of $({\mathcal{M}}(S)^{\lambda},\iota)$ and conjugacy classes of involutions in ${\widehat{\Gamma}}^{\lambda}\cdot\langle\iota\rangle$. By passing to the inverse limit over all such level structures, since $\cap_{{\lambda}\in\Lambda}{\widehat{\Gamma}}^{\lambda}\cdot\langle\iota\rangle=\langle\iota\rangle$, we see that the fixed-point locus ${\mathbb{M}}(S)^{\iota}$ is connected and smooth (and so also irreducible). The proposition then follows from Lemma 3.7. ∎
The closed boundary strata of $\overline{{\mathbb{M}}}(S)$ are parameterized in a ${\check{\Gamma}}^{\pm}(S)$-equivariant way by the simplices of the procongruence curve complex ${\check{C}}(S)$ and, for ${\sigma}\in C(S)\subset{\check{C}}(S)$, there holds $\partial\overline{{\mathbb{M}}}(S)_{\sigma}\cong\overline{{\mathbb{M}}}(S\smallsetminus{\sigma})$, where we denote by $\partial\overline{{\mathbb{M}}}(S)_{\sigma}$ the closed stratum of $\partial\overline{{\mathbb{M}}}(S)=\overline{{\mathbb{M}}}(S)\smallsetminus{\mathbb{M}}(S)$ parameterized by a simplex ${\sigma}\in{\check{C}}(S)$.
For an antiholomorphic involution $\iota$, let us denote by ${\check{C}}(S)^{\iota}_{k}$ the set of $k$-simplices of the simplicial profinite complex ${\check{C}}(S)$ fixed by $\iota$. From Proposition 3.3 and Proposition 3.8, it then follows:
Proposition 3.9.
For $g(S)\leq 2$, the closed irreducible strata of the DM boundary of the fixed-point locus $\overline{{\mathbb{M}}}(S)^{\iota}$ of codimension $k+1$ are parameterized by the fixed-point set ${\check{C}}(S)_{k}^{\iota}$, for all $k\geq 0$. Moreover, ${\check{C}}(S)_{k}^{\iota}$ is the closure of the fixed-point set $C(S)_{k}^{\iota}$ in the profinite set ${\check{C}}(S)_{k}$.
Proof.
Since the closed irreducible strata of the DM boundary $\partial\overline{{\mathbb{M}}}(S)$ of codimension $k+1$ are ${\check{\Gamma}}^{\pm}(S)$-equivariantly parameterized by ${\check{C}}(S)_{k}$, the closed irreducible stratum $\partial\overline{{\mathbb{M}}}(S)^{\iota}_{\sigma}$, parameterized by ${\sigma}\in{\check{C}}(S)_{k}$, is preserved by $\iota$ if and only if ${\sigma}\in{\check{C}}(S)_{k}^{\iota}$ and so, in particular, $\iota\in{\check{\Gamma}}^{\pm}(S)_{\sigma}$. As we observed above, $\partial\overline{{\mathbb{M}}}(S)_{\sigma}\cong\overline{{\mathbb{M}}}(S\smallsetminus{\sigma})$ and ${\check{\Gamma}}^{\pm}(S)_{\sigma}$ acts on it through its quotient ${\check{\Gamma}}^{\pm}(S\smallsetminus{\sigma})$. The image of $\iota$ in ${\Gamma}^{\pm}(S\smallsetminus{\sigma})$ is then also a antiholomorphic involution of this group, which already implies that $\partial\overline{{\mathbb{M}}}(S)_{\sigma}^{\iota}\cong\overline{{\mathbb{M}}}(S\smallsetminus{\sigma})^{\iota}$ is nonempty. By Proposition 3.8, the fixed-point locus is also irreducible. This implies the first claim of the proposition. The last claim follows from the fact that the boundary $\partial\overline{{\mathbb{M}}}(S)^{\iota}=\overline{{\mathbb{M}}}(S)^{\iota}\smallsetminus{{\mathbb{M}}}(S)^{\iota}$ is the closure of the boundary $\partial\overline{{\mathcal{T}}}(S)^{\iota}$ in the boundary $\partial\overline{{\mathbb{M}}}(S)$, which also follows from Proposition 3.8. ∎
3.6. Exceptional automorphisms of ${\Gamma}^{\pm}(S)$ in low genus
Lemma 3.10.
For $S=S_{1,1},S_{1,2}$ or $S_{2}$, there holds:
$H^{1}({\Gamma}^{\pm}(S)/Z({\Gamma}(S)),\mathbb{Z}/2)=\langle\xi\rangle\oplus\langle\zeta\rangle,$
where the epimorphism $\zeta\colon\thinspace{\Gamma}^{\pm}(S)/Z({\Gamma}(S))\to\mathbb{Z}/2$ is induced by the orientation character of ${\Gamma}^{\pm}(S)$ and the restriction of $\xi\colon\thinspace{\Gamma}^{\pm}(S)/Z({\Gamma}(S))\to\mathbb{Z}/2$ to ${\Gamma}(S)/Z({\Gamma}(S))$ is nontrivial.
Proof.
The center $Z({\Gamma}(S))=Z({\Gamma}^{\pm}(S))$ is generated by the hyperelliptic involution $\upsilon$} and the quotient ${\Gamma}(S)_{/{\upsilon}}:={\Gamma}(S)/Z({\Gamma}(S))$ (resp. ${\Gamma}(S)^{\pm}_{/{\upsilon}}:={\Gamma}(S)^{\pm}/Z({\Gamma}(S))$) identifies with the (resp. extended) mapping class group of the quotient surface $S/\langle{\upsilon}\rangle$ marked by the set $B$ of branch points of the natural map $S\to S/\langle{\upsilon}\rangle$. Note that, for $S=S_{1,1},S_{1,2}$ and $S_{2}$, we have $|B|=3,4$ and $6$, respectively. There are then natural epimorphisms ${\Gamma}(S)_{/{\upsilon}}\to\Sigma_{B}$ and ${\Gamma}(S)^{\pm}_{/{\upsilon}}\to\Sigma_{B}$, where $\Sigma_{B}$ is the symmetric group on the set of points $B$.
The inflation-restriction exact sequence associated to the natural short exact sequence $1\to{\Gamma}(S)_{/{\upsilon}}\to{\Gamma}(S)^{\pm}_{/{\upsilon}}\to\mathbb{Z}/2\to 1$ determines the exact sequence:
$0\to\langle\zeta\rangle\to H^{1}({\Gamma}(S)^{\pm}_{/{\upsilon}}),\mathbb{Z}/2)\to H^{1}({\Gamma}(S)_{/{\upsilon}},\mathbb{Z}/2).$
From the results in [13, Section 5.1.3] (essentially, because ${\Gamma}(S)_{/{\upsilon}}$ is normally generated by a Dehn twist), it follows that $H^{1}({\Gamma}(S)_{/{\upsilon}},\mathbb{Z}/2)=\langle\xi\rangle$. In order to prove the lemma, we then only need to show that the cohomology group $H^{1}({\Gamma}(S)^{\pm}_{/{\upsilon}}),\mathbb{Z}/2)$ contains an element whose restriction to ${\Gamma}(S)_{/{\upsilon}}$ is nontrivial.
For this, let us observe that the epimorphisms ${\Gamma}(S)^{\pm}_{/{\upsilon}}\to\Sigma_{B}$ and ${\Gamma}(S)_{/{\upsilon}}\to\Sigma_{B}$ induce on first cohomology groups the monomorphisms $H^{1}(\Sigma_{B},\mathbb{Z}/2)\hookrightarrow H^{1}({\Gamma}(S)^{\pm}_{/{\upsilon}}),\mathbb{Z}/2)$ and $H^{1}(\Sigma_{B},\mathbb{Z}/2)\hookrightarrow H^{1}({\Gamma}(S)_{/{\upsilon}}),\mathbb{Z}/2)$. Since $H^{1}(\Sigma_{B},\mathbb{Z}/2)\cong\mathbb{Z}/2$, the conclusion follows. ∎
By Proposition 3.6, there holds $Z({\widehat{\Gamma}}^{\pm}(S))=Z({\widehat{\Gamma}}(S))=Z({\Gamma}(S))=\langle{\upsilon}\rangle$, for $S=S_{1,1},S_{1,2}$ or $S_{2}$. Hence, with the notations of Lemma 3.10, we have that, for $S=S_{1,1},S_{1,2}$ or $S_{2}$:
(11) $H^{1}({\widehat{\Gamma}}^{\pm}(S)/Z({\widehat{\Gamma}}(S)),Z({\widehat{\Gamma}}(S)))\cong H^{1}({\Gamma}^{\pm}(S)/Z({\Gamma}(S)),Z({\Gamma}(S)))=\langle\xi\rangle\oplus\langle\zeta\rangle,$
where the restriction of the cocycle $\xi$ to ${\widehat{\Gamma}}(S)/Z({\Gamma}(S))$ is nontrivial while the restriction of the cocycle $\zeta$ to ${\widehat{\Gamma}}(S)/Z({\Gamma}(S))$ is trivial.
The epimorphisms $\xi$ and $\zeta$ then determine two automorphisms $\exp\xi$ and $\exp\zeta$ of ${\widehat{\Gamma}}^{\pm}(S)$ (resp. ${\Gamma}^{\pm}(S)$), defined by the assignments $x\mapsto x\cdot\xi(\bar{x})$ and $x\mapsto x\cdot\zeta(\bar{x})$, where $\bar{x}$ denotes the image of $x$ in the quotient ${\widehat{\Gamma}}^{\pm}(S)/Z({\widehat{\Gamma}}(S))$ (resp. ${\Gamma}^{\pm}(S)/Z({\Gamma}(S))$) (cf. the proof of [9, Lemma 3.5]). Note that, by definition, both these automorphisms are not inner.
For the case $S=S_{1,2}$, we then have:
Proposition 3.11.
For $S=S_{1,2}$, there are natural isomorphisms:
  1. (i)
    $\operatorname{Aut}({\widehat{\Gamma}}^{\pm}(S))\cong\operatorname{Aut}(\mathrm{P}{\widehat{\Gamma}}^{\pm}(S))\times\langle\exp\xi\rangle\times\langle\exp\zeta\rangle$;
  2. (ii)
    $\operatorname{Aut}^{\mathbb{I}}({\widehat{\Gamma}}^{\pm}(S))\cong\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\widehat{\Gamma}}^{\pm}(S))\times\langle\exp\zeta\rangle$.
Proof.
We have that ${\widehat{\Gamma}}^{\pm}(S)=\mathrm{P}{\widehat{\Gamma}}^{\pm}(S)\times\langle{\upsilon}\rangle$, because $\upsilon$} determines a splitting of the short exact sequence $1\to\mathrm{P}{\widehat{\Gamma}}^{\pm}(S)\to{\widehat{\Gamma}}^{\pm}(S)\to\mathbb{Z}/2\to 1$.
By the Wells exact sequence associated to the extension $1\to Z({\widehat{\Gamma}}(S))\to{\widehat{\Gamma}}(S)\to\mathrm{P}{\widehat{\Gamma}}^{\pm}(S)\to 1$ and the above identity, since $\operatorname{Aut}(Z({\widehat{\Gamma}}(S)))=\operatorname{Aut}(\mathbb{Z}/2)=\{1\}$, there is a split short exact sequence (cf. [9, Lemma 7.4]):
$0\to H^{1}({\widehat{\Gamma}}^{\pm}(S)/Z({\widehat{\Gamma}}(S)),Z({\widehat{\Gamma}}(S)))\to\operatorname{Aut}({\widehat{\Gamma}}^{\pm}(S))\to\operatorname{Aut}(\mathrm{P}{\widehat{\Gamma}}^{\pm}(S))\to 1.$
The first item of the proposition then follows from (11).
Since the quotient ${\widehat{\Gamma}}(S)/Z({\widehat{\Gamma}}(S))$ is topologically normally generated by the image of a nonseparating Dehn twist $\tau_{\gamma}$ of ${\Gamma}(S)$, there holds $\xi(\bar{\tau}_{\gamma})={\upsilon}$, so that $\exp\xi\notin\operatorname{Aut}^{\mathbb{I}}({\widehat{\Gamma}}^{\pm}(S))$. Instead, since the restriction of $\zeta$ to ${\widehat{\Gamma}}(S)/Z({\widehat{\Gamma}}(S))$ is trivial, we have $\exp\zeta\in\operatorname{Aut}^{\mathbb{I}}({\widehat{\Gamma}}^{\pm}(S))$. The second item of the proposition then also follows. ∎
In the topological case, for $S=S_{1,1},S_{1,2}$ or $S_{2}$, from the isomorphisms (2), Lemma 3.10 and [9, Lemma 3.5], it follows that we have the isomorphisms (cf. also [31, Theorem 1] for the case $S=S_{2}$):
(12) $\operatorname{Aut}({\Gamma}^{\pm}(S))\cong\operatorname{Aut}^{\mathbb{I}}({\Gamma}^{\pm}(S))\times\langle\exp\xi\rangle\hskip 14.22636pt\mbox{and}\hskip 14.22636pt\operatorname{Aut}^{\mathbb{I}}({\Gamma}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}({\Gamma}(S))\times\langle\exp\zeta\rangle.$
@@ 4.1. Preliminary lemmas
Before we proceed to the proof of Theorem 1.4, we need to prove first a series of lemmas. The following definition will be also useful:
Definition 4.1.
We say that a subgroup of ${\check{\Gamma}}^{\pm}(S)$ is $\mathbb{I}$-characteristic if it is preserved by all elements of $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$.
@@
For $S=S_{g,n}$ such that either $g\geq 1$ and $(g,n)\neq(1,2)$ or $g=0$ and $n\geq 5$, there is an epimorphism $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\to\Sigma_{n}$, such that its composition with the homomorphism $\operatorname{inn}\colon\thinspace{\check{\Gamma}}^{\pm}(S)\to\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}(S))$ is the natural epimorphism ${\check{\Gamma}}^{\pm}(S)\to\Sigma_{n}$.
@@
We will first deal with the case $g\geq 1$ and $(g,n)\neq(1,2)$. We will then explain how to adapt the proof to the case $g=0$ and $n\geq 5$.
Let us fix a set of labels $B$ for the punctures on $S$. For a $1$-simplex $\{{\gamma}_{0},{\gamma}_{1}\}\in C(S)$ such that the two curves ${\gamma}_{0},{\gamma}_{1}$ bound a $1$-punctured annulus on $S$, we define the marked topological type of $\{{\gamma}_{0},{\gamma}_{1}\}$ as the data of the topological type of $\{{\gamma}_{0},{\gamma}_{1}\}$ plus the label $b\in B$ which marks the puncture on the annulus with boundary ${\gamma}_{0}\cup{\gamma}_{1}$. Note that the hypothesis $S\neq S_{1,2}$ makes sure that this is well defined.
There are $n$ marked topological types of such simplices and they are in bijective correspondence with the $\mathrm{P}\Gamma(S)$-orbits of $1$-simplices $\{{\gamma}_{0},{\gamma}_{1}\}$ of the given topological type. The action of ${\Gamma}^{\pm}(S)$ on $C(S)$ then permutes these orbits and the natural epimorphism ${\Gamma}^{\pm}(S)\to\Sigma_{n}$ can be recovered from this action.
For a $1$-simplex $\{{\gamma}_{0},{\gamma}_{1}\}\in{\check{C}}(S)$ with the topological type of a $1$-punctured annulus on $S$ described above, we then define its marked topological type as the marked topological type of any $1$-simplex $\{{\gamma}_{0}^{\prime},{\gamma}_{1}^{\prime}\}\in C(S)$ in the $\mathrm{P}{\check{\Gamma}}(S)$-orbit of $\{{\gamma}_{0},{\gamma}_{1}\}$. Since there holds ${\check{C}}(S)_{1}/\mathrm{P}{\check{\Gamma}}(S)=C(S)_{1}/\mathrm{P}\Gamma(S)$, this is well defined.
By [9, Theorem 5.5], the natural representation $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}({\check{C}}(S))$ preserves the topological types of simplices in ${\check{C}}(S)$. This determines an action on marked topological types of $1$-simplices ${\sigma}\in{\check{C}}(S)$ with the topological type of a $1$-punctured annulus on $S$ if and only if, ${\sigma}^{\prime}=f\cdot{\sigma}$, for some $f\in\mathrm{P}{\check{\Gamma}}(S)$, implies that, for every $\phi\in\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$, there exists an $f^{\prime}\in\mathrm{P}{\check{\Gamma}}(S)$ such that $\phi\cdot{\sigma}^{\prime}=f^{\prime}\cdot\phi\cdot{\sigma}$.
The pure mapping class group $\mathrm{P}\Gamma(S)$ is the subgroup of the extended mapping class group ${\Gamma}^{\pm}(S)$ generated by Dehn twists. Hence, $\mathrm{P}{\check{\Gamma}}(S)$ is an $\mathbb{I}$-characteristic subgroup of ${\check{\Gamma}}^{\pm}(S)$ in the sense of Definition 4.1. This implies that $\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}(S))$ is a normal subgroup of $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$. Therefore, the natural representation $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\to\operatorname{Aut}({\check{C}}(S))$ satisfies the property stated above. The action of $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$ on marked topological types of $1$-simplices of the topological type of a $1$-punctured annulus then determines a representation $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\to\Sigma_{n}$ with the desired properties.
For $g=0$ and $n\geq 5$, we consider instead $1$-simplices $\{{\gamma}_{0},{\gamma}_{1}\}\in C(S)$ such that the curves ${\gamma}_{0},{\gamma}_{1}$ together bound a $1$-punctured annulus and, moreover, one of the two curves bounds a $2$-punctured disc on $S$. This is a well defined topological type of $1$-simplices for $n\geq 5$ which we then mark by labeling the puncture on the annulus with boundary ${\gamma}_{0}\cup{\gamma}_{1}$. We can then proceed exactly as in the case $g\geq 1$ and $(g,n)\neq(1,2)$ treated above. ∎
@@
The group $\mathrm{P}{\check{\Gamma}}(S)$ is an $\mathbb{I}$-characteristic subgroup of $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$ and, for $d(S)>1$ and $S\neq S_{1,2}$, $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$ is an $\mathbb{I}$-characteristic subgroup of ${\check{\Gamma}}^{\pm}(S)$.
@@
The pure mapping class group $\mathrm{P}\Gamma(S)$ is the subgroup of the extended mapping class group ${\Gamma}^{\pm}(S)$ generated by Dehn twists. Hence, $\mathrm{P}{\check{\Gamma}}(S)$ is an $\mathbb{I}$-characteristic subgroup of both ${\check{\Gamma}}^{\pm}(S)$ and $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$.
In order to prove that, for $d(S)>1$ and $S\neq S_{1,2}$, $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$ is an $\mathbb{I}$-characteristic subgroup of ${\check{\Gamma}}^{\pm}(S)$, it is enough to show that $\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ is a normal subgroup of $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$. This immediately follows from the identity $\operatorname{Inn}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))=\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))\cap\ker\Theta$, where $\Theta\colon\thinspace\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\to\Sigma_{n}$ is the representation constructed in Lemma 4.2. ∎
Remark 4.4.
Lemma 4.3 fails for $S=S_{1,2}$. In fact, the exceptional automorphism $\exp\zeta\in\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$ constructed in Section 3.6 has the property that $\exp\zeta(\mathrm{P}{\check{\Gamma}}^{\pm}(S))\neq\mathrm{P}{\check{\Gamma}}^{\pm}(S)$.
From Lemma 4.3, it follows that the elements of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))$ are compatible with the augmentation map $\mathrm{P}{\check{\Gamma}}^{\pm}(S)\to\mathbb{Z}/2$ and so preserve the set of antiholomorphic involutions. Moreover, by (ii) of Proposition 3.5, for $g(S)\leq 2$, the sets of conjugacy classes of antiholomorphic involutions in $\mathrm{P}\Gamma^{\pm}(S)$ and $\mathrm{P}{\widehat{\Gamma}}^{\pm}(S)$ can be identified. We have:
@@
For $g(S)=0$, there is only one ${\widehat{\Gamma}}(S)^{\pm}$-conjugacy class of antiholomorphic involutions in $\mathrm{P}{\widehat{\Gamma}}(S)^{\pm}$, on which $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\widehat{\Gamma}}(S)^{\pm})$ acts naturally.
@@
Antiholomorphic involutions in $\mathrm{P}\Gamma(S)^{\pm}$ do not swap the punctures of $S$. Therefore, all the punctures lie in $\operatorname{Fix}(\iota)$, which, in particular, is not empty. Since $g(S)=0$, $\operatorname{Fix}(\iota)$ is also connected and separating. Hence, an antiholomorphic involution in $\mathrm{P}\Gamma(S)^{\pm}$ is determined by the cyclic order of the punctures on $\operatorname{Fix}(\iota)$. Since ${\Gamma}(S)^{\pm}$ acts transitively on such orderings, any two antiholomorphic involutions in $\mathrm{P}\Gamma(S)^{\pm}$ are conjugated by an element of ${\Gamma}(S)^{\pm}$. The conclusion then follows from (ii) of Proposition 3.5. ∎